OpenStudy (s3a):

Question: If A is a 3 by 5 matrix, what information do you have about the nullspace of A? Answer: N(A) has dimension at least 2 and at most 5. My trouble: How do we know this?

OpenStudy (anonymous):

m>=r m=3 so 0=<r=<3 dim N(A) = n-r=5-r 2<=5-r<=5

OpenStudy (s3a):

You're using the notation where m is the number of rows and r is the rank, right?

OpenStudy (anonymous):

yes m=#rows n=#columns r=rank

OpenStudy (s3a):

So m >= r always holds for the nullspace of a matrix?

OpenStudy (anonymous):

rank is #pivot rows right? can it be more than #of rows?

OpenStudy (anonymous):

it is just a fact for every matrices

OpenStudy (anonymous):

i mean you can't have more pivot 'rows' than rows

OpenStudy (s3a):

No, it cannot. I now get the answer to that sub-question.

OpenStudy (s3a):

I mean my sub-question not part (b). Let me look at 3(b) again

OpenStudy (s3a):

Why is dim N(A) = n-r=5-r?

OpenStudy (s3a):

Also is dim N(A) = dim ("the x variable vector in Ax = 0")?

OpenStudy (anonymous):

i don't know whether you've heard the mit ocw lecture or not, but you should know dim N(A) = n -r

OpenStudy (anonymous):

matrix is 3by 5 so n=5

OpenStudy (s3a):

I haven't. Is N(A) = n - r something I am supposed to "just memorize"?

OpenStudy (s3a):

dim N(A) = n - r that is

OpenStudy (anonymous):

well dim N(A) is number of vectors in basis of nullspace of A

OpenStudy (anonymous):

think about the basis of nullspace of A think how you get those bases

OpenStudy (s3a):

(where n and r are of the matrix A) (Sorry for the potentially dumb question, I'm kind of braindead at the moment but, like I said, I cannot take a break.)

OpenStudy (s3a):

(My break will be when I sleep.)

OpenStudy (anonymous):

can you think how you obtain the null space of A?

OpenStudy (anonymous):

then it is clear cut that dim N(A) = n-r

OpenStudy (s3a):

you find the vector x in Ax = 0

OpenStudy (s3a):

?

OpenStudy (anonymous):

yeah i'm asking do you know the exact procedures

OpenStudy (s3a):

get A^(-1) and multiply both sides?

OpenStudy (s3a):

(on the left)

OpenStudy (s3a):

i mean left multiplication on each side of the equation

OpenStudy (anonymous):

since there are r pivot columns, there are n-r free columns and there are n-r special solutions to the null space. And they form basis for nullspace. so dim N(A)=n-r

OpenStudy (anonymous):

you can do that only when a is invertible

OpenStudy (anonymous):

you use elimination process to compute nullspace right?

OpenStudy (anonymous):

why don't you solve some examples to get dimN(A)=n-r i think you know the definitions

OpenStudy (s3a):

when a is invertible, like you said. I'm not grasping something fundamental though: what is the nullspace of A? is it x in Ax = 0? In other words, what object's dimension is n - r?

OpenStudy (anonymous):

nullspace(A) is a subspace in R^n, which a vector in it satisfies Ax=0

OpenStudy (anonymous):

it is just whole solutions to Ax=0

OpenStudy (anonymous):

solutions form a subspace, so they are called null'space'

OpenStudy (s3a):

ok so it's a dimension of a vector space in which a vector x makes Ax = 0?

OpenStudy (anonymous):

for example, when A is invertible, x is only 0 and the N(A)={0}

OpenStudy (anonymous):

no nullspace is just a name of subspace

OpenStudy (anonymous):

dim N(A) is what you are saying

OpenStudy (s3a):

ya for two seconds, my brain was thinking about dim N(A).

OpenStudy (s3a):

Will you be here in 1..75 hours?

OpenStudy (anonymous):

you need to carry on the elimination yourself to see dim N(A) = n-r

OpenStudy (s3a):

I have to go eat now.

OpenStudy (anonymous):

i'm sorry

OpenStudy (s3a):

(If you keep writing, I will come back and read what you said.)

OpenStudy (anonymous):

i have to go :(

OpenStudy (s3a):

o :(

OpenStudy (anonymous):

i have a pdf about nullspace. do you want it?

OpenStudy (anonymous):

http://ocw.mit.edu/courses/mathematics/18-06sc-linear-algebra-fall-2011/ax-b-and-the-four-subspaces/solving-ax-0-pivot-variables-special-solutions/ download the lecture summary for that lecture. it has the just the right information for you. good luck

OpenStudy (s3a):

thanks and sry for leaving abruptbly